OpenStudy (didierr):

Estimate 24% of 200.

OpenStudy (didierr):

5,000 50 5 0.5

OpenStudy (anonymous):

What is 1 % of 200?

OpenStudy (anonymous):

Take that number and multiply it with 24 :)

OpenStudy (anonymous):

48

OpenStudy (didierr):

i got 4800

OpenStudy (anonymous):

\[1\%=\frac{200}{100}=?\] \[24*?=??\]

OpenStudy (anonymous):

if 24% of 100 is 24 just double it to get 24% of 200 so 24%*2=48%

OpenStudy (futureboy1):

48

OpenStudy (futureboy1):

trust me my answer is correct

OpenStudy (anonymous):

Thats true ash_heather. But for future questions where it isnt that easy, you need to know the general way to solve it.

OpenStudy (anonymous):

yea i know

OpenStudy (didierr):

its 5000

OpenStudy (anonymous):

How did you get 5000?

OpenStudy (futureboy1):

i just ask my trusted computer intelligence

OpenStudy (anonymous):

@didierr lol no its not check it on google just put in 24% of 200

OpenStudy (futureboy1):

but to solve the problem by hand is very easy

OpenStudy (anonymous):

\[1\%=\frac{200}{100}=2\] So: \[24\%=2*24=48\]

OpenStudy (futureboy1):

i use cortana

OpenStudy (futureboy1):

its a windows 10 thing

OpenStudy (anonymous):

@Tommynaiter thats right :)

OpenStudy (futureboy1):

sorry i meant 48%